# Fast Robust Kernel Regression through Sign Gradient Descent with Early Stopping

### v6

@@ A.2 Kernel Gradient Descent and Kernel Gradient Flow
whose solution is stated in Lemma 5.
Lemma 5.
The solution to the differential equation in Equation 19 is
$\begin{bmatrix}\bm{\hat{f}}(t)\\ \bm{\hat{f}^{*}}(t)\end{bmatrix}=\begin{bmatrix}\bm{I}\\ \bm{K^{*}}\bm{K}^{-1}\end{bmatrix}(\bm{I}-\exp(-t\bm{K}))\bm{y}$
@@ A.3 Kernel Regression with the $\ell_{\infty}$ and $\ell_{1}$ Norms
In contrast to ridge regression, where the formulations in $\bm{\alpha}$ and $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ are equivalent in the sense that the latter can always be obtained from the first through multiplication with $[\bm{K}^{\top},\bm{K^{*}}{{}^{\top}}]^{\top}$, the solutions of Equations 20 and 21 cannot be obtained from the solutions of Equations 13 and 14.
@@
Finally, for completeness, we present the analog of Proposition 4 for
$[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ in Proposition 6.
Proposition 6.
@@
$\frac{\partial\bm{f^{{}^{+}}}(t)}{\partial t}=-\text{sign}\left(\frac{\partial}{\partial\bm{f^{{}^{+}}}(t)}\left(\left\|\bm{y^{{}^{+}}}-\bm{f^{{}^{+}}}(t)\right\|_{\bm{K^{**}}}^{2}\right)\right),\ \bm{f^{{}^{+}}}(0)=\bm{0}.$
@@
Appendix B Additional Experiments
@@
In this section, we extend the results of Tables 4 and 5 by repeating the experiments for the four kernels in Table 6. Since, at the time of writing, the KQRT-T implementation only supported the Gaussian kernel, the algorithm is omitted in this section. The results are presented in Tables 7 to 14 and agree with those of Section 6.2.
@@
Table 7: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for $\nu=0.5$ (Laplace kernel). The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 8: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for $\nu=0.5$ (Laplace kernel). The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Table 9: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for $\nu=1.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 10: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for $\nu=1.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Table 11: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for $\nu=2.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 12: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for $\nu=2.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Table 13: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for the Cauchy kernel. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 14: The 2.5th, 50th and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for the Cauchy kernel. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Appendix C Proofs
Proof of Lemma 1.
@@ Appendix C Proofs
Solving the differential equations in the same way as in Lemma 2, we obtain
@@
Proof of Lemma 3.
$\left(\bm{K}+\lambda\bm{I}\right)^{-1}$ $=\left(\bm{K}+\lambda\bm{I}\right)^{-1}\underbrace{(\bm{K}+\lambda\bm{I}-\lambda\bm{I})\bm{K}^{-1}}_{=\bm{I}}=\left(\bm{I}-\lambda\left(\bm{K}+\lambda\bm{I}\right)^{-1}\right)\bm{K}^{-1}$
$=\left(\bm{I}-\left(\bm{I}+1/\lambda\cdot\bm{K}\right)^{-1}\right)\bm{K}^{-1}.$
Proof of Proposition 1.
By using the fact that $\bm{K}$, $\bm{K}^{-1}$ and $\bm{\Delta_{K}}:=\left(\left(\bm{I}+t\bm{K}\right)^{-1}-\exp(-t\bm{K})\right)$ commute, we obtain
@@
$\mathbb{E}(\bm{\hat{\alpha}}_{\text{KRR}}(1/t))-\bm{\alpha_{0}}=\left(\bm{K}+1/t\cdot\bm{I}\right)^{-1}\mathbb{E}(\bm{y})-\bm{\alpha_{0}}=\left(\bm{K}+1/t\cdot\bm{I}\right)^{-1}\bm{K}\bm{\alpha_{0}}-\bm{\alpha_{0}}$
@@
$=\text{Tr}\left(\left(\bm{\Sigma_{\alpha}}\cdot 1/t^{2}\cdot\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}+\sigma_{\varepsilon}^{2}\cdot 1.2985^{2}\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}\right)\bm{k}(\bm{x^{*}})\bm{k}(\bm{x^{*}})^{\top}\right)$
$\leq 1.6862\cdot\text{Tr}\left(\left(\bm{\Sigma_{\alpha}}\cdot 1/t^{2}\cdot\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}+\sigma_{\varepsilon}^{2}\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}\right)\bm{k}(\bm{x^{*}})\bm{k}(\bm{x^{*}})^{\top}\right)$
@@
$\text{sign}\left(-\frac{\partial}{\partial\bm{\beta}(t)}\left(\left\|\bm{y}-\bm{X}\bm{\beta}(t)\right\|_{2}^{2}\right)\right)=\text{sign}\left(\bm{X}^{\top}\bm{y}-\bm{X}^{\top}\bm{X}\bm{\beta}(t)\right)$
$=\text{sign}\left(\begin{bmatrix}\vrule\\ (\bm{X}^{\top}\bm{y})_{i}-s_{ii}\beta_{i}(t)\\ \vrule\end{bmatrix}\right)=\begin{bmatrix}\vrule\\ \text{sign}\left((\bm{X}^{\top}\bm{y})_{i}-s_{ii}\beta_{i}(t)\right)\\ \vrule\end{bmatrix}.$
Since element $i$ in the vector only depends on $\bm{\beta}$ through $\beta_{i}$,
$\frac{\partial\bm{\beta}(t)}{\partial t}$ $=\text{sign}\left(-\frac{\partial}{\partial\bm{\beta}(t)}\left(\left\|\bm{y}-\bm{X}\bm{\beta}(t)\right\|_{2}^{2}\right)\right)$
$\iff\frac{\partial\beta_{i}(t)}{\partial t}$ $=\text{sign}((\bm{X}^{\top}\bm{y})_{i}-s_{ii}\beta_{i}(t))$
=$=\begin{cases}1&\text{if }(\bm{X}^{\top}\bm{y})_{i}>s_{ii}\beta(t)\\ -1&\text{if }(\bm{X}^{\top}\bm{y})_{i}<s_{ii}\beta(t)\\ 0&\text{if }(\bm{X}^{\top}\bm{y})_{i}=s_{ii}\beta(t)\\ \end{cases},\ i=1,2,\dots p.$
Thus, with $\beta_{i}(0)=0$, $\beta_{i}(t)=\pm t$, depending on the sign of $(\bm{X}^{\top}\bm{y})_{i}/s_{ii}$. Once $\beta_{i}=(\bm{X}^{\top}\bm{y})_{i}/s_{ii}$ (and $\text{sign}(s_{ii}\beta_{i}-(\bm{X}^{\top}\bm{y})_{i})=0$), then $\beta_{i}=(\bm{X}^{\top}\bm{y})_{i}/s_{ii}$. That is,
@@
$\text{sign}\left(-\frac{\partial}{\partial\bm{\alpha}(t)}\left(\left\|\bm{y}-\bm{K}\bm{\alpha}(t)\right\|_{\bm{K}^{-1}}^{2}\right)\right)=\text{sign}\left(\bm{y}-\bm{K}\bm{\alpha}(t)\right)$
$=\text{sign}\left(\begin{bmatrix}\vrule\\ y_{i}-k_{ii}\alpha_{i}(t)\\ \vrule\end{bmatrix}\right)=\begin{bmatrix}\vrule\\ \text{sign}(y_{i}-k_{ii}\alpha_{i}(t))\\ \vrule\end{bmatrix}.$
Since element $i$ in the vector only depends on $\bm{\alpha}$ through $\alpha_{i}$,
$\frac{\partial\bm{\alpha}(t)}{\partial t}=\text{sign}\left(-\frac{\partial}{\partial\bm{\alpha}(t)}\left(\left\|\bm{y}-\bm{K}\bm{\alpha}(t)\right\|_{\bm{K}^{-1}}^{2}\right)\right)$
$\iff$ $\frac{\partial\alpha_{i}(t)}{\partial t}=\text{sign}(y_{i}-k_{ii}\alpha_{i}(t))=\begin{cases}1&\text{if }y_{i}>k_{ii}\alpha_{i}(t)\\ -1&\text{if }y_{i}<k_{ii}\alpha_{i}(t)\\ 0&\text{if }y_{i}=k_{ii}\alpha_{i}(t)\\ \end{cases},\ i=1,2,\dots n.$
Thus, with $\alpha_{i}(0)=0$, $\alpha_{i}(t)=\pm t$, depending on the sign of $y_{i}/k_{ii}$. Once $\alpha_{i}=y_{i}/k_{ii}$ (and $\text{sign}(y_{i}-k_{ii}\alpha_{i})=0$), then $\alpha_{i}=y_{i}/k_{ii}$. That is,

### v7

@@ A.2 Kernel Gradient Descent and Kernel Gradient Flow
whose solution is given by Equation 19.
$\begin{bmatrix}\bm{\hat{f}}(t)\\ \bm{\hat{f}^{*}}(t)\end{bmatrix}=\begin{bmatrix}\bm{I}\\ \bm{K^{*}}\bm{K}^{-1}\end{bmatrix}(\bm{I}-\exp(-t\bm{K}))\bm{y}.$ (19)
@@ A.3 Kernel Regression with the $\ell_{\infty}$ and $\ell_{1}$ Norms
In contrast to ridge regression, where the formulations in $\bm{\alpha}$ and $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ are equivalent in the sense that the latter can always be obtained from the first through multiplication with $[\bm{K}^{\top},\bm{K^{*}}{{}^{\top}}]^{\top}$, the solutions of Equations 20 and 21 cannot be obtained from the solutions of Equations 12 and 13.
@@
Finally, for completeness, we present the analog of Proposition 5.1 for
$[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ in Proposition A.2.
Proposition A.2.
@@
$\frac{\partial\bm{\hat{f}^{{}^{+}}}(t)}{\partial t}=-\text{sign}\left(\frac{\partial}{\partial\bm{\hat{f}^{{}^{+}}}(t)}\left(\left\|\bm{y^{{}^{+}}}-\bm{\hat{f}^{{}^{+}}}(t)\right\|_{\bm{K^{**}}}^{2}\right)\right),\ \bm{\hat{f}^{{}^{+}}}(0)=\bm{0}.$
@@
Appendix B Computationally Efficient Cross-Validation
For models of the type $\bm{\hat{f}}=\bm{H}\bm{y}$, where $\bm{H}\in\mathbb{R}^{n\times n}$ is a matrix that does not depend on $\bm{y}$, the fast leave-one-out cross-validation, LOOCV, and generalized cross-validation, GCV, statistics are defined according to
$\text{LOOCV}:=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{y_{i}-\hat{f}(\bm{x_{i}})}{1-H_{ii}}\right)^{2}$
$\text{GCV}:=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{y_{i}-\hat{f}(\bm{x_{i}})}{1-\text{Tr}(\bm{H})/n}\right)^{2}.$
These two methods tend to be faster than k-fold cross-validation and provide similar performance. For KRR, where $\bm{\hat{f}}=\bm{K}(\bm{K}+\lambda\bm{I})^{-1}\bm{y}$, we trivially obtain $\bm{H}=\bm{K}(\bm{K}+\lambda\bm{I})^{-1}$. For other penalties than the ridge penalty, it is however not trivial to obtain $\bm{H}$. Tibshirani, (1996) noted that the lasso objective, $\frac{1}{2}\left\|\bm{y}-\bm{X}\bm{\beta}\right\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}$, can be reformulated as $\frac{1}{2}\left\|\bm{y}-\bm{X}\bm{\beta}\right\|_{2}^{2}+\frac{\lambda}{2}\bm{\beta}^{\top}\bm{B}\bm{\beta}$, where $\bm{B}\in\mathbb{R}^{p\times p}$ is a diagonal matrix with $B_{ii}=2/|\beta_{i}|$. If $\bm{B}$ is treated as a constant, this formulation becomes a weighted ridge regression problem which is minimized by $\bm{\hat{\beta}}=\bm{X}(\bm{X}^{\top}\bm{X}+\lambda\bm{B})^{-1}\bm{X}^{\top}\bm{y}$, and thus, for a given $\bm{\hat{\beta}}$, the author proposed approximating $\bm{H}$ as $\bm{X}(\bm{X}^{\top}\bm{X}+\lambda\bm{\hat{B}})^{-1}\bm{X}^{\top}$, where $\bm{\hat{B}}$ is a diagonal matrix with $\hat{B}_{ii}=2/|\bm{\hat{\beta}}|$.
Applying these ideas in the kernel formulation, we obtain, for an $\bm{A}$ such that $\frac{1}{2}\bm{\alpha}^{\top}\bm{A}\bm{\alpha}=\|\bm{\alpha}\|$,
$\bm{\hat{\alpha}}$ $=\operatorname*{arg\,min}_{\bm{\alpha}\in\mathbb{R}^{n}}\frac{1}{2}\left\|\bm{y}-\bm{K}\bm{\alpha}\right\|_{\bm{K}^{-1}}^{2}+\lambda\|\bm{\alpha}\|$
$=\operatorname*{arg\,min}_{\bm{\alpha}\in\mathbb{R}^{n}}\frac{1}{2}\left\|\bm{y}-\bm{K}\bm{\alpha}\right\|_{\bm{K}^{-1}}^{2}+\frac{\lambda}{2}\bm{\alpha}^{\top}\bm{A}\bm{\alpha}=(\bm{K}+\lambda\bm{A})^{-1}\bm{y}.$
Thus $\bm{H}$ can be approximated by $\bm{K}(\bm{K}+\lambda\bm{\hat{A}})^{-1}$, where $\bm{\hat{A}}\in\mathbb{R}^{n\times n}$ is a matrix such that $\frac{1}{2}\bm{\hat{\alpha}}^{\top}\bm{\hat{A}}\bm{\hat{\alpha}}=\|\bm{\hat{\alpha}}\|$. For the $\ell_{1}$ norm, $\|\bm{\hat{\alpha}}\|_{1}=\sum_{i=1}^{n}|\hat{\alpha}_{i}|$, this is obtained if $\bm{\hat{A}}$ is a diagonal matrix with $\hat{A}_{ii}=2/|\hat{\alpha}_{i}|$. For the $\ell_{\infty}$ norm, $\|\bm{\hat{\alpha}}\|_{\infty}=\max_{i}|\hat{\alpha}_{i}|$, it is obtained if $\bm{\hat{A}}$ is a matrix with only zeros, except for $\hat{A}_{mm}=2/|\hat{\alpha}_{m}|$, where $m=\operatorname*{arg\,max}_{i}|\hat{\alpha}_{i}|$.
In Tables 6 and 7, we present the analogs of Tables 4 and 5 for LOOCV and GCV. As expected, the two methods perform faster than 10-fold cross-validation. However, for $\ell_{\infty}$, the performance in terms of test $R^{2}$ is much lower than for 10-fold cross-validation. For KRR, LOOCV performs well in terms of test $R^{2}$.
Table 6: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for $\ell_{\infty}$ in combination with LOOCV and GCV on the different data sets, without amplified outliers. Compared to 10-fold cross validation LOOCV and GCV perform faster, but the cost of decreased predictive performance, especially for $\ell_{\infty}$.
Data Method Type of Cross- Validation Computation Time [s] 50%, (2.5%, 97.5%) Test $R^{2}$ 50%, (2.5%, 97.5%)
Airfoil Sound Pressure $\ell_{\infty}$ LOOCV $49.1,\ (48.2,57.2)$ $-0.01,\ (-0.38,0.04)$
GCV $50.0,\ (48.7,57.9)$ $-0.01,\ (-0.37,0.15)$
KRR LOOCV $1.21,\ (0.891,3.11)$ $0.55,\ (-0.29,0.84)$
GCV $1.43,\ (0.910,3.35)$ $0.33,\ (-4.74,0.75)$
California House Values $\ell_{\infty}$ LOOCV $43.1,\ (41.2,50.3)$ $-0.02,\ (-0.36,0.05)$
GCV $43.5,\ (41.7,52.8)$ $-0.03,\ (-0.37,0.00)$
KRR LOOCV $1.26,\ (0.928,2.25)$ $0.60,\ (-0.13,0.88)$
GCV $1.26,\ (0.841,2.16)$ $0.47,\ (-0.39,0.83)$
Steel Energy Consumption $\ell_{\infty}$ LOOCV $54.1,\ (51.7,63.1)$ $0.00,\ (-0.38,0.71)$
GCV $54.3,\ (51.5,68.5)$ $0.57,\ (-0.08,0.98)$
KRR LOOCV $0.891,\ (0.793,1.63)$ $0.99,\ (0.97,1.00)$
GCV $0.892,\ (0.797,1.57)$ $0.98,\ (0.71,1.00)$
Superconductor Critical Temperature $\ell_{\infty}$ LOOCV $52.4,\ (51.2,56.0)$ $-0.00,\ (-0.42,0.25)$
GCV $52.8,\ (51.6,56.1)$ $0.44,\ (0.01,0.74)$
KRR LOOCV $1.79,\ (1.02,2.91)$ $0.69,\ (0.12,0.87)$
GCV $2.30,\ (1.15,3.83)$ $0.46,\ (-2.71,0.79)$
U.K. Temperature $\ell_{\infty}$ LOOCV $48.5,\ (46.6,64.7)$ $-0.05,\ (-0.38,0.03)$
GCV $48.6,\ (46.8,61.3)$ $-0.03,\ (-0.38,0.10)$
KRR LOOCV $1.01,\ (0.860,1.59)$ $0.38,\ (-0.30,0.68)$
GCV $0.993,\ (0.796,1.65)$ $0.28,\ (-6.81,0.71)$
Table 7: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for $\ell_{\infty}$ in combination with LOOCV and GCV on the different data sets, with amplified outliers. Compared to 10-fold cross validation LOOCV and GCV perform faster, but the cost of decreased predictive performance, especially for $\ell_{\infty}$.
Data Method Type of Cross- Validation Computation Time [s] 50%, (2.5%, 97.5%) Test $R^{2}$ 50%, (2.5%, 97.5%)
Airfoil Sound Pressure $\ell_{\infty}$ LOOCV $52.8,\ (48.9,60.4)$ $-0.02,\ (-14.57,0.01)$
GCV $56.1,\ (49.7,64.9)$ $-0.02,\ (-14.33,0.08)$
KRR LOOCV $0.840,\ (0.760,0.925)$ $0.30,\ (-1.12,0.66)$
GCV $0.844,\ (0.754,0.997)$ $-0.05,\ (-189.78,0.69)$
California House Values $\ell_{\infty}$ LOOCV $44.7,\ (42.4,50.7)$ $-0.02,\ (-0.21,0.01)$
GCV $45.0,\ (43.0,50.3)$ $-0.03,\ (-1.28,0.29)$
KRR LOOCV $0.985,\ (0.832,3.58)$ $0.29,\ (-0.62,0.77)$
GCV $1.01,\ (0.833,4.69)$ $0.16,\ (-37.61,0.72)$
Steel Energy Consumption $\ell_{\infty}$ LOOCV $54.5,\ (52.3,68.9)$ $-0.04,\ (-0.94,0.12)$
GCV $55.1,\ (52.3,65.7)$ $0.01,\ (-0.97,0.88)$
KRR LOOCV $0.869,\ (0.753,1.04)$ $0.79,\ (-0.80,0.97)$
GCV $0.889,\ (0.762,1.05)$ $0.13,\ (-405.97,0.96)$
Superconductor Critical Temperature $\ell_{\infty}$ LOOCV $52.9,\ (51.3,61.0)$ $-0.04,\ (-0.90,0.02)$
GCV $53.3,\ (51.5,62.1)$ $0.47,\ (-0.19,0.78)$
KRR LOOCV $1.04,\ (0.993,1.23)$ $0.54,\ (-0.46,0.75)$
GCV $1.06,\ (0.987,1.29)$ $0.13,\ (-37.16,0.62)$
U.K. Temperature $\ell_{\infty}$ LOOCV $51.2,\ (48.2,59.1)$ $-0.05,\ (-4.55,0.00)$
GCV $51.2,\ (48.3,58.7)$ $-0.05,\ (-1.12,0.00)$
KRR LOOCV $0.906,\ (0.818,1.13)$ $0.18,\ (-2.03,0.56)$
GCV $0.887,\ (0.817,1.14)$ $-0.01,\ (-61.69,0.58)$
Appendix C Additional Experiments
@@
In this section, we extend the results of Tables 4 and 5 by repeating the experiments for the four kernels in Table 8. Since, at the time of writing, the KQRT-T implementation only supported the Gaussian kernel, this algorithm is omitted in this section. The results are presented in Tables 9 to 16 and agree with those of Section 6.2.
@@
Table 9: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for $\nu=0.5$ (Laplace kernel). The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 10: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for $\nu=0.5$ (Laplace kernel). The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Table 11: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for $\nu=1.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 12: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for $\nu=1.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Table 13: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for $\nu=2.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 14: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for $\nu=2.5$. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Table 15: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, without amplified outliers, for the Cauchy kernel. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster.
@@
Table 16: The 2.5th, 50th, and 97.5th percentiles of computation time and test $R^{2}$ for the different methods and data sets, with amplified outliers, for the Cauchy kernel. The five robust methods perform very similarly in terms of test $R^{2}$, while KSGD performs one to two orders of magnitude faster. The two non-robust methods, KGD/KRR, do not perform as well as the robust methods.
@@
Appendix D Proofs
Proof of Proposition 3.1.
@@ Appendix D Proofs
Solving the differential equations in the same way as for Equation 10, we obtain
@@
Proof of Proposition 4.1.
By using the fact that $\bm{K}$, $\bm{K}^{-1}$, and $\bm{\Delta_{K}}:=\left(\left(\bm{I}+t\bm{K}\right)^{-1}-\exp(-t\bm{K})\right)$ commute, we obtain
@@
$\mathbb{E}(\bm{\hat{\alpha}}_{\text{KRR}}(1/t))-\bm{\alpha_{0}}$
$=\left(\bm{K}+1/t\cdot\bm{I}\right)^{-1}\mathbb{E}(\bm{y})-\bm{\alpha_{0}}=\left(\bm{K}+1/t\cdot\bm{I}\right)^{-1}\bm{K}\bm{\alpha_{0}}-\bm{\alpha_{0}}$
@@
$=$ $\text{Tr}\left(\left(\bm{\Sigma_{\alpha}}\cdot 1/t^{2}\cdot\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}\right.\right.$
$\left.\left.+\sigma_{\varepsilon}^{2}\cdot 1.2985^{2}\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}\right)\bm{k}(\bm{x^{*}})\bm{k}(\bm{x^{*}})^{\top}\right)$
$\leq$ $1.6862\cdot\text{Tr}\left(\left(\bm{\Sigma_{\alpha}}\cdot 1/t^{2}\cdot\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}\right.\right.$
$\left.\left.+\sigma_{\varepsilon}^{2}\left(1/t\cdot\bm{I}+\bm{K}\right)^{-2}\right)\bm{k}(\bm{x^{*}})\bm{k}(\bm{x^{*}})^{\top}\right)$
@@
$\text{sign}\left(-\frac{\partial}{\partial\bm{\hat{\beta}}(t)}\left(\left\|\bm{y}-\bm{X}\bm{\hat{\beta}}(t)\right\|_{2}^{2}\right)\right)=\text{sign}\left(\bm{X}^{\top}\bm{y}-\bm{X}^{\top}\bm{X}\bm{\hat{\beta}}(t)\right)$
$=\text{sign}\left(\begin{bmatrix}\vrule\\ (\bm{X}^{\top}\bm{y})_{i}-s_{ii}\hat{\beta}_{i}(t)\\ \vrule\end{bmatrix}\right)=\begin{bmatrix}\vrule\\ \text{sign}\left((\bm{X}^{\top}\bm{y})_{i}-s_{ii}\hat{\beta}_{i}(t)\right)\\ \vrule\end{bmatrix}.$
Since element $i$ in the vector only depends on $\bm{\hat{\beta}}$ through $\hat{\beta}_{i}$,
$\frac{\partial\bm{\hat{\beta}}(t)}{\partial t}$ $=\text{sign}\left(-\frac{\partial}{\partial\bm{\hat{\beta}}(t)}\left(\left\|\bm{y}-\bm{X}\bm{\hat{\beta}}(t)\right\|_{2}^{2}\right)\right)$
$\iff\frac{\partial\hat{\beta}_{i}(t)}{\partial t}$ $=\text{sign}((\bm{X}^{\top}\bm{y})_{i}-s_{ii}\hat{\beta}_{i}(t))$
=$=\begin{cases}1&\text{if }(\bm{X}^{\top}\bm{y})_{i}>s_{ii}\hat{\beta}(t)\\ -1&\text{if }(\bm{X}^{\top}\bm{y})_{i}<s_{ii}\hat{\beta}(t)\\ 0&\text{if }(\bm{X}^{\top}\bm{y})_{i}=s_{ii}\hat{\beta}(t)\\ \end{cases},\ i=1,2,\dots p.$
Thus, with $\hat{\beta}_{i}(0)=0$, $\hat{\beta}_{i}(t)=\pm t$, depending on the sign of $(\bm{X}^{\top}\bm{y})_{i}/s_{ii}$. Once $\hat{\beta}_{i}=(\bm{X}^{\top}\bm{y})_{i}/s_{ii}$ (and $\text{sign}(s_{ii}\hat{\beta}_{i}-(\bm{X}^{\top}\bm{y})_{i})=0$), then $\hat{\beta}_{i}=(\bm{X}^{\top}\bm{y})_{i}/s_{ii}$. That is,
@@
$\text{sign}\left(-\frac{\partial}{\partial\bm{\hat{\alpha}}(t)}\left(\left\|\bm{y}-\bm{K}\bm{\hat{\alpha}}(t)\right\|_{\bm{K}^{-1}}^{2}\right)\right)=\text{sign}\left(\bm{y}-\bm{K}\bm{\hat{\alpha}}(t)\right)$
$=\text{sign}\left(\begin{bmatrix}\vrule\\ y_{i}-k_{ii}\hat{\alpha}_{i}(t)\\ \vrule\end{bmatrix}\right)=\begin{bmatrix}\vrule\\ \text{sign}(y_{i}-k_{ii}\hat{\alpha}_{i}(t))\\ \vrule\end{bmatrix}.$
Since element $i$ in the vector only depends on $\bm{\hat{\alpha}}$ through $\hat{\alpha}_{i}$,
$\frac{\partial\bm{\hat{\alpha}}(t)}{\partial t}=\text{sign}\left(-\frac{\partial}{\partial\bm{\hat{\alpha}}(t)}\left(\left\|\bm{y}-\bm{K}\bm{\hat{\alpha}}(t)\right\|_{\bm{K}^{-1}}^{2}\right)\right)$
$\iff$ $\frac{\partial\hat{\alpha}_{i}(t)}{\partial t}=\text{sign}(y_{i}-k_{ii}\hat{\alpha}_{i}(t))=\begin{cases}1&\text{if }y_{i}>k_{ii}\hat{\alpha}_{i}(t)\\ -1&\text{if }y_{i}<k_{ii}\hat{\alpha}_{i}(t)\\ 0&\text{if }y_{i}=k_{ii}\hat{\alpha}_{i}(t)\\ \end{cases},\ i=1,2,\dots n.$
Thus, with $\hat{\alpha}_{i}(0)=0$, $\hat{\alpha}_{i}(t)=\pm t$, depending on the sign of $y_{i}/k_{ii}$. Once $\hat{\alpha}_{i}=y_{i}/k_{ii}$ (and $\text{sign}(y_{i}-k_{ii}\hat{\alpha}_{i})=0$), then $\hat{\alpha}_{i}=y_{i}/k_{ii}$. That is,